LABORATORY DETECTION AND PURE ROTATIONAL SPECTRUM OF VCl^{+} BY VELOCITY MODU-LATION SPECTROSCOPY

\underline{D. T. HALFEN} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of VCl^{+} has been measured using millimeter-wave direct absorption techniques incorporating velocity modulation for ion selection. This work is the first laboratory observation of this species. VCl^{+} was created in an AC glow discharge of gas-phase VCl_{4} and argon. Rotational transitions of the ^{35}Cl and ^{37}Cl isotopomers have been measured originating in four fine structure components of this ion, each which exhibit hyperfine structure arising from the 51V nuclear spin (I = 7/2). The current data suggest a ^{4}Φ_{r} or ^{4}Δ_{r} ground state. Additional spectroscopic measurements are in progress and the latest results will be presented.